# J-HERTz: J-PLUS Heritage Exploration of Radio Targets at z 5

### v1

@@ 4.3 Stellar masses and star formation rates of galaxies
To explore host galaxy properties, we divide the sample into three groups based on their distribution in Fig. 12: radio-loud galaxies, radio-quiet galaxies with blue IR colors, and radio-quiet galaxies with red IR colors. Figure 13 shows the relationship between radio luminosity, stellar mass and sSFR for these groups. Radio-quiet galaxies follow a stellar mass–radio luminosity relation similar to the radio-IR correlation, as the IR traces the older stellar population that dominates the stellar mass when no AGN is present. On the other hand, the radio luminosity in radio-loud galaxies is independent of the stellar mass (Hardcastle2023). Both radio-quiet galaxy populations have similar sSFR values. As mentioned above, red W1-W2 colors observed in some radio-quiet galaxies may result from nuclear activity in a spatially resolved galaxy that was not classified as a QSO by BANNJOS, although some of these sources also reach high sSFR values consistent with starbursts (% have sSFR ).
@@ 4.4 Stranger things
Finally, there are many objects classified as stars with associated radio emission that could be coming from other nearby or undetected objects. A search within the general J-PLUS catalog identifies 6,391 stars exhibiting extended radio emission with a nearby object that is not classified as a star. We perform some filtering criteria explained in Section 3.3 to find the most reliable radio star candidates, which we flag with radio_star_candidate = ‘reliable’, while the rest have ‘unreliable’.
@@
5 Conclusions and future work
@@
Most galaxies are radio-quiet and lie in the IR–radio relation (left panel of Fig. 10). The majority of radio-quiet galaxies with blue W1-W2 colors are passive galaxies, showing intermediate masses () and low sSFRs (0.01-0.25 ). Radio-quiet galaxies with red W1-W2 colors have tentatively lower masses () and highest sSFRs (0.04-0.34 ), reaching values compatible with starburst galaxies (% of them have values larger than 0.5 ). This suggests that the red IR colors are associated with ongoing starbursts activity. However, around half of galaxies with red IR colors (W1-W2 0.25) and available SDSS spectroscopy are classified as QSOs by Almeida2023 and might be AGN with a resolved host.
@@
Notably, there are 14,000 radio-loud galaxies (at r 21) with no clear optical or IR signs of nuclear activity. Most of these galaxies have extended radio emission at low frequencies likely tracing past jet activity. These are also the most massive () and most quenched galaxies (1 spread of the sSFR of 0.01-0.09 , Fig. 14). About half of them are compatible with low excitation radio galaxies (LERGs) according to the spectral classification provided by Drake2024, while we estimate that 20% could be optically quiescent radio galaxies with residual or absent nuclear activity, and extended relic radio lobes at low frequencies being the only evidence of the past activity. Further studies of their radio morphologies, spectral indices, surface brightness and core prominences are necessary to completely characterize these sources.
@@
The J-HERTz catalog is available as part of the J-PLUS database. In the near future, the catalog will be updated with the upcoming J-PLUS DR4, adding 2,000 with improved photo-z estimates (Hernán-Caballero in prep.) We will use J-HERTz to select a clean sample of star-forming galaxies and compare SFRs and radio luminosities for this population, to characterize the radio-SFR correlation (Fernández Gil et al. in prep). This will allow us to determine SFRs of radio detected galaxies missing precise optical photometry and also establish a selection criteria between star-forming and AGN radio galaxies. Additional radio surveys at higher frequencies such as the Arecibo Legacy Fast ALFA H I (ALFALFA; Haynes2018) or the first data release of the APERture Tile In Focus array (Apertif; vanCappellen2022) will trace the HI emission in these sources. Finally, we aim to further characterize the host galactic properties, estimate the black hole accretion rates using emission line ratios thanks to the precise narrow band photometry of J-PLUS, and compare them with past AGN activity seen in older and extended jets detected with low-frequency radio.
Catalog parameters
| Parameter name | Description | Units | ||
| TILE_ID | Tile id of the reference band image of the object in J-PLUS. | |||
| NUMBER | Number id assigned by SExtractor to the object in J-PLUS. | |||
| Source_name | Object Identifier (ILT name) in LoTSS. | |||
| RA | Radio right ascension from LoTSS. | deg | ||
| DEC | Radio declination from LoTSS. | deg | ||
| Opt_RA | Right ascension of the optical/IR counterpart from WISE/SDSS. | deg | ||
| Opt_DEC | Declination of the optical/IR counterpart from WISE/SDSS. | deg | ||
| ang_Dist | Angular distance between the optical/IR counterpart and the J-PLUS matching object. | arcsec | ||
| Total_flux | 144-MHz total flux density from LoTSS. | mJy | ||
| E_Total_flux | Error on total flux density from LoTSS. | mJy | ||
| LAS | Radio angular size estimate from LoTSS. | arcsec | ||
| Resolved | Resolved flag from LoTSS. | |||
| CLASS_GALAXY_mean | Mean probability of being a galaxy by BANNJOS. | |||
| CLASS_QSO_mean | Mean probability of being a QSO by BANNJOS. | |||
| CLASS_STAR_mean | Mean probability of being a star by BANNJOS. | |||
| CLASS_GALAXY_std | Standard deviation of probability of being a galaxy by BANNJOS. | |||
| CLASS_QSO_std | Standard deviation of probability of being a QSO by BANNJOS. | |||
| CLASS_STAR_std | Standard deviation of probability of being a star by BANNJOS. | |||
| LoTSS_zphot | Photo-z estimate from LoTSS. | |||
| LoTSS_zphot_err | Photo-z error from LoTSS. | |||
| flag_qual | Predicted photo-z quality of LoTSS, 0 if bad, 1 if good. | |||
| JPLUS_zphot | Photo-z estimate from J-PLUS. | |||
| JPLUS_zphot_err | Photo-z error from J-PLUS. | |||
| ODDS | Integral of P(z) within JPLUS_zphot ± 0.03*(1+JPLUS_zphot). | |||
| Combined_zphot | Combined photo-z from J-PLUS and LoTSS. | |||
| Combined_zphot_err | Combined photo-z error. | |||
| zspec | Spectroscopic redshift from SDSS, DESI or HETDEX. | |||
| zspec_err | Spectroscopic redshift error when available. | |||
| z_best | Best available redshift. | |||
| z_best_err | Best available redshift error. | |||
| z_source | Source of the best available redshift. | |||
| Size | Radio linear size obtained from LAS using z_best for the cosmology. | kpc | ||
| L_144 | Radio luminosity obtained from Total_flux using z_best for the cosmology. | W/Hz | ||
| Radio_loudness | Radio-to-IR ratio as | dex | ||
| FLUX_AUTO_Lambda | AUTO flux densities for the 12 J-PLUS bands. | |||
| FLUX_RELERR_AUTO_Lambda | Relative errors of the AUTO flux density values for the 12 J-PLUS bands. | |||
| FLAGS | Sextractor flags for the 12 J-PLUS bands. | |||
| MASK_FLAGS | Mask flags for the 12 J-PLUS bands. | |||
| mag_w1 | Magnitude in WISE band 1 (taken from LoTSS). | mag | ||
| magerr_w1 | Magnitude error in WISE band 1, or blank for upper limit. | mag | ||
| mag_w2 | Magnitude in WISE band 2 (taken from LoTSS). | mag | ||
| magerr_w2 | Magnitude error in WISE band 2, or blank for upper limit. | mag | ||
| mag_w3 | Magnitude in WISE band 3 (taken from LoTSS). | mag | ||
| magerr_w3 | Magnitude error in WISE band 3, or blank for upper limit. | mag | ||
| mag_w4 | Magnitude in WISE band 4 (taken from LoTSS). | mag | ||
| magerr_w4 | Magnitude error in WISE band 4, or blank for upper limit. | mag | ||
| Stellar_mass | Stellar mass from CIGALE for z_best. | |||
| Stellar_mass_err | Stellar mass error from CIGALE for z_best. | |||
| SFR | SFR from CIGALE for z_best. | /yr | ||
| SFR_err | SFR_err from CIGALE for z_best. | /yr | ||
| SFR10Myrs | SFR10Myrs from CIGALE for z_best. | /yr | ||
| SFR10Myrs_err | SFR10Myrs_err from CIGALE for z_best. | /yr | ||
| SFR100Myrs | SFR100Myrs from CIGALE for z_best. | /yr | ||
| SFR100Myrs_err | SFR100Myrs_err from CIGALE for z_best. | /yr | ||
| close_objects | Flag indicating an object closer than 5 arcsec. | |||
| 0 if not present, 1 if present. | ||||
| incompatible_spec_class | Flag indicating a SDSS/DESI spectroscopic class different from the BANNJOS classification. | |||
| 0 if no, 1 if yes. | ||||
| unreliable_spec_z | Flag indicating an unreliable SDSS/DESI spectroscopic redshift. | |||
| 0 if no, 1 if yes. | ||||
| candidate_radio_star | Flag indicating a potential radio star candidate, with values ‘reliable’ or ‘unreliable’ | |||
| ambiguous_classification | Flag indicating an object with several BANNJOS classifications, | |||
| mostly due to several detections in GAIA. 0 if not, 1 if yes. |

### v2

@@ IV.3 Stellar masses and star formation rates of galaxies
To explore host galaxy properties, we divide the sample into three groups based on their distribution in Fig. 12: radio-loud galaxies, radio-quiet galaxies with blue IR colors, and radio-quiet galaxies with red IR colors. Figure 13 shows the relationship between radio luminosity, stellar mass and sSFR for these groups. Radio-quiet galaxies follow a stellar mass–radio luminosity relation similar to the radio-IR correlation, as the IR traces the older stellar population that dominates the stellar mass when no AGN is present. On the other hand, the radio luminosity in radio-loud galaxies is independent of the stellar mass (Hardcastle et al., 2023). Both radio-quiet galaxy populations have similar sSFR values. As mentioned above, red W1-W2 colors observed in some radio-quiet galaxies may result from nuclear activity in a spatially resolved galaxy that was not classified as a QSO by BANNJOS, although some of these sources also reach high sSFR values consistent with starbursts (% have sSFR ).
@@ IV.4 Stranger things
Finally, there are many objects classified as stars with associated radio emission that could be coming from other nearby or undetected objects. A search within the general J-PLUS catalog identifies 6,391 stars exhibiting extended radio emission with a nearby object that is not classified as a star. We perform some filtering criteria explained in Section III.3 to find the most reliable radio star candidates, which we flag with radio_star_candidate = ‘reliable’, while the rest have ‘unreliable’.
@@
V Conclusions and future work
@@
Most galaxies are radio-quiet and lie in the IR–radio relation (left panel of Fig. 10). The majority of radio-quiet galaxies with blue W1-W2 colors are passive galaxies, showing intermediate masses () and low sSFRs (0.01-0.25 ). Radio-quiet galaxies with red W1-W2 colors have tentatively lower masses () and highest sSFRs (0.04-0.34 ), reaching values compatible with starburst galaxies (% of them have values larger than 0.5 ). This suggests that the red IR colors are associated with ongoing starbursts activity. However, around half of galaxies with red IR colors (W1-W2 0.25) and available SDSS spectroscopy are classified as QSOs by Almeida et al. (2023) and might be AGN with a resolved host.
@@
Notably, there are 14,000 radio-loud galaxies (at r 21) with no clear optical or IR signs of nuclear activity. Most of these galaxies have extended radio emission at low frequencies likely tracing past jet activity. These are also the most massive () and most quenched galaxies (1 spread of the sSFR of 0.01-0.09 , Fig. 14). About half of them are compatible with low excitation radio galaxies (LERGs) according to the spectral classification provided by Drake et al. (2024), while we estimate that 20% could be optically quiescent radio galaxies with residual or absent nuclear activity, and extended relic radio lobes at low frequencies being the only evidence of the past activity. Further studies of their radio morphologies, spectral indices, surface brightness and core prominences are necessary to completely characterize these sources.
@@
The J-HERTz catalog is available as part of the J-PLUS database. In the near future, the catalog will be updated with the upcoming J-PLUS DR4, adding 2,000 with improved photo-z estimates (Hernán-Caballero in prep.) We will use J-HERTz to select a clean sample of star-forming galaxies and compare SFRs and radio luminosities for this population, to characterize the radio-SFR correlation (Fernández Gil et al. in prep). This will allow us to determine SFRs of radio detected galaxies missing precise optical photometry and also establish a selection criteria between star-forming and AGN radio galaxies. Additional radio surveys at higher frequencies such as the Arecibo Legacy Fast ALFA H I (ALFALFA; Haynes et al., 2018) or the first data release of the APERture Tile In Focus array (Apertif; van Cappellen et al., 2022) will trace the HI emission in these sources. Finally, we aim to further characterize the host galactic properties, estimate the black hole accretion rates using emission line ratios thanks to the precise narrow band photometry of J-PLUS, and compare them with past AGN activity seen in older and extended jets detected with low-frequency radio.